6. Draw the conjugate acid of phenetidine and explain why it is highly soluble in water. Make sure to reference any relevant intermolecular forces as part of your answer.

6b. The solubility of dulcin in water is 1.21 g/L at room temperature. Assuming that you added a total of 82.5 mL of water between the reaction setup and the recrystallization process, how much dulcin (in mg) would remain in solution after recrystallization if you allowed the flask to cool to room temperature?
Answer #1

6. a.

Structure of conjugate acid of phenetidine is given below.

NH3

The conjugate acid is ionic, hence hydrophilic .

The intermolecular force present between the conjugate acid and water is ion-dipole force. Due to this force conjugate acid of phenetidine is highly soluble in water .

b.

Solute is dulcin.

solubility = \small \frac{mass \: of \: solute (g) }{volume\, of \: water(L)}     =   1.21  \tiny \frac{g}{L}

given , Volume of water = 82.5 mL = 0.0825 L

now, \small \frac{mass \: of \: solute (g) }{0.0825 (L)} = 1.21  \tiny \frac{g}{L}

or, mass of solute = 1.21 ( \tiny \frac{g}{L} ) \small \times 0.0825 (L) = 0.0998 g = 0.0998 g \small \times\tiny \frac{1000 \, mg}{1\, g}  = 99.825 mg .

Therefore mass of dulcin ( solute) would remain in solution after recrystallization at room temperature is 99.825 mg.